OpenStudy (anonymous):

Find the height of a triangle with area of 42 square centimeters and a base of 7cm.

OpenStudy (vishweshshrimali5):

12

OpenStudy (anonymous):

28

OpenStudy (anonymous):

@sarsar i solved a similar prob. and you could have tried it by urself..

OpenStudy (diyadiya):

1/2bh=42 1/2*7*h=42 h=42*2/7 h=12

OpenStudy (anonymous):

i dont feel good i have 10 minutes left in my exam i just need help ! ):

OpenStudy (anonymous):

12 is the right i did a mistake